# Let’s Influence Algorithms Together: How Millions of Fans Build Collective Understanding of Algorithms and Organize Coordinated Algorithmic Actions

### v1

@@ Abstract.
Previous research pays attention to how users strategically understand and consciously interact with algorithms but mainly focuses on an individual level, making it difficult to explore how users within communities could develop a collective understanding of algorithms and organize collective algorithmic actions. Through a two-year ethnography of online fan activities, this study investigates 43 core fans who always organize large-scale fans’ collective actions and their corresponding general fan groups. This study aims to reveal how these core fans mobilize millions of general fans through collective algorithmic actions. These core fans reported the rhetorical strategies used to persuade general fans, the steps taken to build a collective understanding of algorithms, and the collaborative processes that adapt collective actions across platforms and cultures. Our findings highlight the key factors that enable computer-supported collective algorithmic actions and extend collective action research into large-scale domain targeting algorithms.
@@ 1. Introduction
Past HCI research has explored how various kinds of users strategically develop an understanding of platform algorithms and subsequently negotiate with these algorithms (DeVito et al., 2017; Cotter, 2019; Karizat et al., 2021; DeVos et al., 2022). For instance, Cotter discussed how Instagram influencers develop an instrumental interaction of algorithms to increase their visibility by discerning the operational principles of these algorithms (Cotter, 2019). Yu et al. discovered that delivery workers formed micro-scale algorithmic communities by establishing private social media groups for mutual support. In these groups, they collaboratively developed a folk understanding of algorithms and shared insights into algorithmic functioning and resistance strategies (Yu et al., 2022). These studies have investigated how individuals or small groups of individuals make use of or resist algorithms and progressively develop their understanding of algorithms. However, at the collective level, particularly when considering larger groups of millions or even tens of millions, it remains unclear to HCI scholars how communities collectively understand algorithms and how collective actions towards algorithms could be developed and organized.
@@
In contrast to the academic community’s lack of research on large-scale collective action targeting algorithms, such behaviors are on the rise and are having a significant impact on both platforms and society (Bonini and Treré, 2024; Treré and Bonini, 2024). Citizens are increasingly dissatisfied with their passive role in the face of algorithms and are seeking to enhance their influence through algorithmic activism (Velkova and Kaun, 2021; Bonini and Treré, 2024). For instance, political groups have previously engaged in Google bombing by utilizing predictions of Google’s algorithms to develop strategies for designing websites, thereby influencing Google search outcomes (Gillespie, 2019). In 2018, activists manipulated Google’s algorithm and organized a large group of political allies on Reddit, massively upvoting Trump-related posts, which caused Google’s ranking algorithm to associate the term ‘idiot’ with his image (Haynes, 2018; Gillespie, 2019).
@@
As previous algorithm scholars have pointed out, a significant trend in digital platforms is that more and more groups are using algorithms as tools for social movements and for gaining visibility (Bonini and Treré, 2024; Treré and Bonini, 2024; Gillespie, 2019). These groups amplify their influence over algorithms and hijack platform algorithms to achieve their own goals (Treré and Bonini, 2024). However, there is a lack of research in academia on how these groups mobilize and organize large-scale actions targeting algorithms.
Our research uses large-scale (millions) fan-driven, computer-supported collective actions targeting algorithms (occurring across multiple countries and platforms) as a case study to expand the understanding of HCI and related scholars regarding such large-scale collective algorithmic actions.
@@
Our work’s key contributions to the HCI community are as follows:
Through a two-year ethnographic study, complemented by semi-structured interviews and focus groups, we conducted an in-depth analysis of how millions of fans orchestrate influential collective algorithmic actions.
We advanced folk theories of algorithms from an individual dimension to a collective understanding and explored how these folk theories evolve into well-defined, context-specific algorithmic action. We also extend computer-supported collective action studies to the field of algorithms.
We explored the potential for large-scale, computer-supported collective action by examining fan-driven algorithmic efforts capable of mobilizing millions of users. Our study, encompassing multiple fan communities, nations, and platforms, demonstrates not only the broad applicability of our theory to future large-scale collective actions but also potential strategies for overcoming large-scale collective action dilemmas.
@@ 2.1. Folk Theories of Algorithm
HCI studies have examined how users perceive various aspects of algorithms, such as how they affect user visibility on social media (Alvarado and Waern, 2018; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022), the processes behind content moderation (Mayworm et al., 2024; Moran et al., 2022), how algorithms facilitate user matching on dating apps and how news is curated by algorithms (Ngo and Krämer, 2022) These previous studies reveal how users build and refine their understanding of algorithms based on their day-to-day experiences and observations. In this process of testing and updating, users can develop algorithmic folk theories that are sufficiently intricate and flexible (Karizat et al., 2021). As they interact with algorithms and refine their understanding through ongoing experiences and observations, these folk theories become more nuanced. Users adjust their beliefs based on new information and changing interactions, allowing their folk theories to evolve through an iterative process in response to varying algorithmic behaviors and outcomes (DeVito, 2021; Zhang et al., 2023).
@@ 2.2. Computer-Supported Collective Action
Shaw et al. (Shaw et al., 2014) proposed five stages in computer-supported collective action:
Stage 1. Identify a problem and others who care about it. Recognize key issues affecting a specific public and find others who share the concern.
Stage 2. Generate, debate, and select viable solutions. Focus on creating and choosing practical solutions once aligned with a committed group.
Stage 3. Coordinate and prepare for action. Plan the necessary steps, assign leadership, define tasks, set milestones, and clarify timelines and responsibilities.
Stage 4. Take action. Maintain coordination to adapt to new information and changing circumstances.
Stage 5. Assess, document, and follow up. Evaluate, document, and review outcomes to prepare for future work efficiently.
However, it’s important to note that many collective actions on social media fail to achieve their goals; for instance, over 99% of Change.org petitions aimed at improving privacy protections never reach ”victory” status (Wu et al., 2022; Huang et al., 2015). To conclude, many of these efforts fail because they bypass key stages in the model proposed by Shaw et al. (Shaw et al., 2014), such as effective issue framing, consensus-building, and the structured coordination of collective actions. Users also often lack the specific computational tools to collectively action.
@@ 2.3. Transformative Fandom
Although these studies have explored the transformative potential of fans and the phenomenon of fans developing collective actions targeting algorithms, research is still lacking on how these large-scale collective actions are organized within fan communities. While earlier studies have provided a theoretical foundation for collective algorithmic imaginaries among fans, our work primarily contributes by examining how these collective algorithmic understandings undergo iterative development through collective efforts, extend to actionable levels, and become integrated into everyday practices across the community.
@@ 3. Method
In this study on fans’ collective algorithmic actions, we define ”core fans” as those who hold organizational power within the fan community, possess significant reputations, and directly influence the organization of collective fan actions. In contrast, ”general fans” are those who participate in collective algorithmic actions but do not directly engage in core decision-making or shape the overall strategy of the actions; their primary role is to execute tasks as directed.
More specifically, we conducted an in-depth two-year ethnographic study from 2021 to 2023, focusing on interactions with 43 core fans in China. The study not only involved observing these 43 Chinese core fans but also included tracking and analyzing the daily activities of their respective fan clubs and the broader general fan groups, all with their consent. To supplement the empirical data, we also conducted one-on-one interviews with 15 general fans from their fan clubs.
The study received approval from the university’s ethics review board. Throughout the research process, the researcher maintained transparency regarding their role and obtained informed consent from all participants. We consistently prioritized listening to the voices of participants, critically reflecting on our positionality as researchers, and conscientiously examining potential biases.
Throughout these two years, we engaged with these Chinese core fans biweekly to gather their reflections on their daily collective algorithmic actions. These reflections encompassed both the internal strategies within the core fan group, as well as their influence on the general fan group. This ethnographic method allowed us to understand the complex behavioral patterns of fans in collective algorithmic actions and how they coordinate across different levels of fan groups to drive successful collective action.
@@ 3.1. Participating Role
Furthermore, these two researchers also engaged in the algorithmic practices typical of general fans, dedicating 1 to 2 hours daily, from Monday to Friday, to perform routine activities such as repetitive liking, reposting, and commenting, and adjusting specific strategies across different platforms. The platforms where our researchers conducted these algorithmic practices included YouTube, Weibo, X (formerly Twitter), Douyin, TikTok, Xiaohongshu, and other popular platforms where fan activities are active. Additionally, they actively joined and observed the online communities of general fans within the fan clubs corresponding to these 43 core fans, participating in their discussions.
3.2. Participants: Core Fans (N=43), General Fans (N=15)
@@
The 43 Chinese core fans (CF1 to CF43) we interviewed all actively participated in organizing daily fan algorithmic actions. In addition, each organized over 30 specialized collective algorithmic actions targeting specific events, such as brand campaigns involving their idol, showcasing their extensive experience in coordinating such activities. The platforms impacted by the algorithmic actions discussed in these interviews included major fan activity platforms like YouTube, Weibo, X (formerly Twitter), Bilibili, Douyin, TikTok, Xiaohongshu, WeChat Channel and various music charts. Among the Chinese core fans we interviewed, 4 (CF1 to CF4) managed fan clubs with 100,000 to 200,000 followers on their social media accounts, 21 (CF5 to CF25) managed clubs with 200,000 to 500,000 followers, 11 (CF26 to CF36) managed clubs with 500,000 to 1,000,000 followers, and 7 (CF37 to CF43) managed clubs with more than 1,000,000 followers. According to our participants, the follower count on a fan club’s social media account is a more accurate indicator of the number of people involved in collective algorithmic actions compared to the follower count on the idol’s own social media account. In their view, following the idol might indicate an appreciation for the idol personally, but it doesn’t necessarily imply active participation in algorithmic actions. In contrast, those who follow a fan club are much more likely to be highly motivated to engage in fan-driven algorithmic activities. Therefore, the data from these fan clubs can assist HCI researchers in understanding the scale of their organized algorithmic actions. These interviewed Chinese core fans could directly participate in and influence their fan clubs’ decisions and judgments regarding algorithmic folk understanding and collective algorithmic actions. Additionally, all the core fans interviewed have their own social media account besides their fan club, while each had over 5,000 followers on social media platforms, with a monthly average readership exceeding 2 million, and their social media content primarily focused on their idol and fan community, thus affirming their core fans status (sufficient influence driven by their own fan identity).

### v2

@@ Abstract.
Previous research pays attention to how users strategically understand and consciously interact with algorithms but mainly focuses on an individual level, making it difficult to explore how users within communities could develop a collective understanding of algorithms and organize collective algorithmic actions. Through a two-year ethnography of online fan activities, this study investigates 43 core fans who always organize large-scale fans collective actions and their corresponding general fan groups. This study aims to reveal how these core fans mobilize millions of general fans through collective algorithmic actions. These core fans reported the rhetorical strategies used to persuade general fans, the steps taken to build a collective understanding of algorithms, and the collaborative processes that adapt collective actions across platforms and cultures. Our findings highlight the key factors that enable computer-supported collective algorithmic actions and extend collective action research into the large-scale domain targeting algorithms.
@@ 1. Introduction
Past HCI research has explored how various kinds of users strategically develop an understanding of platform algorithms and subsequently negotiate with these algorithms (DeVito et al., 2017; Cotter, 2019; Karizat et al., 2021; DeVos et al., 2022). For instance, Cotter discussed how Instagram influencers develop an instrumental interaction of algorithms to increase their visibility by discerning the operational principles of these algorithms (Cotter, 2019). Yu et al. discovered that delivery workers formed micro-scale algorithmic communities by establishing private social media groups for mutual support. In these groups, they collaboratively developed a folk understanding of algorithms and shared insights into algorithmic functioning and resistance strategies (Yu et al., 2022). These studies have investigated how individuals or small groups of individuals make use of or resist algorithms and progressively develop their understanding of algorithms. However, at the collective level, particularly when considering larger groups of millions or even tens of millions, it remains unclear to HCI scholars how communities collectively understand algorithms and how collective actions towards algorithms could be developed and organized. Understanding these large-scale actions is crucial due to their societal and technological implications. They challenge algorithmic integrity, expose gaps in platform governance, and reflect how users actively reshape algorithms to their advantage. Studying these behaviors offers insights for creating fairer, more transparent, and resilient platforms while contributing to discussions on algorithmic accountability and ethics.
@@
In contrast to the academic community’s lack of research on large-scale collective action targeting algorithms, such behaviors are on the rise and are having a significant impact on both platforms and society (Bonini and Treré, 2024; Treré and Bonini, 2024). Citizens are increasingly dissatisfied with their passive role in the face of algorithms and are seeking to enhance their influence through algorithmic activism (Velkova and Kaun, 2021; Bonini and Treré, 2024). For instance, political groups have previously engaged in Google bombing by utilizing predictions of Google’s algorithms to develop strategies for designing websites, thereby influencing Google search outcomes (Gillespie, 2019). In 2018, activists manipulated Google’s algorithm and organized a large group of political allies on Reddit, massively upvoting Trump-related posts, which caused Google’s ranking algorithm to associate the term ‘idiot’ with his image (Haynes, 2018; Gillespie, 2019). Thus, such large-scale collective actions, while empowering for communities, may also pose significant challenges. These actions can undermine platform integrity by distorting algorithmic outcomes, overcrowding spaces meant for public discourse, and eroding trust in platforms. For instance, fan-driven actions (Zhang et al., 2023) can overshadow critical societal discussions, manipulate public perceptions, and promote artificial visibility at the expense of organic user engagement. Platforms must balance accommodating grassroots participation with maintaining the authenticity of their algorithmic outputs, which becomes increasingly difficult as users refine their manipulation tactics.
@@
As previous scholars of critical algorithm studies have pointed out, a significant trend in digital platforms is that more and more groups are using algorithms as tools for social movements and for gaining visibility (Bonini and Treré, 2024; Treré and Bonini, 2024; Gillespie, 2019). These groups amplify their influence over algorithms and hijack platform algorithms to achieve their own goals (Treré and Bonini, 2024). However, there is a lack of research in academia on how these groups mobilize and organize large-scale actions targeting algorithms.
Our research uses large-scale (millions) fan-driven, computer-supported collective actions targeting algorithms (occurring across multiple countries and platforms) as a case study to expand the understanding of HCI and related scholars regarding such large-scale collective algorithmic actions. By studying fan-driven algorithmic actions, we can better understand the mechanisms and motivations behind these behaviors and inform strategies to address their potentially harmful impacts. Our findings can help platforms develop governance models that are both resilient to manipulation and inclusive of diverse user voices. This research also contributes to ongoing discussions about the ethical implications of user-algorithm interactions, providing actionable insights for creating systems that are fairer, more transparent, and resistant to abuse. These contributions are particularly valuable for the CHI community, which focuses on the intersection of human behavior, technology design, and governance (Duvall and Heckemeyer, 2018; Jung, 2012).
@@
Our work’s key contributions to the HCI community are as follows: (1) Through a two-year ethnographic study, complemented by semi-structured interviews and focus groups, we conducted an in-depth analysis of how millions of fans orchestrate influential collective algorithmic actions. (2) We advanced folk theories of algorithms from an individual dimension to a collective understanding and explored how these folk theories evolve into well-defined, context-specific algorithmic action. We also extend computer-supported collective action studies to the field of algorithms. (3) We explored the potential for large-scale, computer-supported collective action by examining fan-driven algorithmic efforts capable of mobilizing millions of users. Our study, encompassing multiple fan communities, nations, and platforms, demonstrates not only the broad applicability of our theory to future large-scale collective actions but also potential strategies for overcoming large-scale collective action dilemmas.
@@ 2.1. Folk Theories of Algorithm
HCI studies have examined how users perceive various aspects of algorithms, such as how they affect user visibility on social media (Alvarado and Waern, 2018; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022), the processes behind content moderation (Mayworm et al., 2024; Moran et al., 2022), how algorithms facilitate user matching on dating apps and how news is curated by algorithms (Ngo and Krämer, 2022). These previous studies reveal how users build and refine their understanding of algorithms based on their day-to-day experiences and observations. In this process of testing and updating, users can develop algorithmic folk theories that are sufficiently intricate and flexible (Karizat et al., 2021). As they interact with algorithms and refine their understanding through ongoing experiences and observations, these folk theories become more nuanced. Users adjust their beliefs based on new information and changing interactions, allowing their folk theories to evolve through an iterative process in response to varying algorithmic behaviors and outcomes (DeVito, 2021; Zhang et al., 2023).
@@ 2.2. Computer-Supported Collective Action
Shaw et al. (Shaw et al., 2014) identified five key stages in computer-supported collective action. The first stage involves identifying a problem and finding others who share an interest in addressing it. The second stage focuses on generating, debating, and selecting viable solutions to the problem. In the third stage, participants coordinate and prepare for action, followed by the fourth stage, where the collective action is executed. Finally, the fifth stage emphasizes assessing, documenting, and following up on the outcomes of the action, ensuring that lessons are learned and future efforts are informed.
However, it is important to note that many collective actions on social media fail to achieve their goals; for instance, over 99% of Change.org petitions aimed at improving privacy protections never reach ”victory” status (Wu et al., 2022; Huang et al., 2015). To conclude, many of these efforts fail because they bypass key stages in the model proposed by Shaw et al. (Shaw et al., 2014), such as effective issue framing, consensus-building, and the structured coordination of collective actions. Users also often lack the specific computational tools to collectively action.
@@ 2.3. Transformative Fandom
Although these studies have explored the transformative potential of fans and the phenomenon of fans developing collective actions targeting algorithms, research is still lacking on how these large-scale collective actions are organized within fan communities. While earlier studies have provided a theoretical foundation for collective algorithmic imaginaries among fans (Zhang et al., 2023; Zheng and Xiao, 2023; Zhang and Negus, 2020), our work primarily contributes by examining how these collective algorithmic understandings undergo iterative development through collective efforts, extend to actionable levels, and become integrated into everyday practices across the community.
@@ 3. Method
In this research, we conducted an in-depth two-year ethnographic study from 2021 to 2023, focusing on interactions with 43 core fans from 12 famous fanclub in China. The study involved observing these 43 core fans, as well as tracking and analyzing the daily activities of their respective fan clubs and the broader general fan groups, all with their consent. As a supplement to the ethnographic observations, we conducted one-on-one interviews with all 43 core fans and organized three focus group discussions to delve deeper into their perspectives. Additionally, to broaden our understanding, we conducted one-on-one interviews with 15 general fans from these fan clubs to gather insights into their experiences and roles within the community.
3.1. Ethnography Context: Define “Core Fans” and “General Fans”
In this study on fans’ collective algorithmic actions, we define ”core fans” as those who hold organizational power within the fan community, possess remarkable reputations, and directly influence the organization of collective fan actions. In contrast, ”general fans” are those who participate in collective algorithmic actions but do not directly engage in core decision-making or shape the overall strategy of the actions; their primary role is to execute tasks as directed.
To be more detailed, in Chinese pop fan communities, interactions between core fans and general fans primarily occur online. Core fans, known for their dedication and critical influence within the fanbase, often go beyond virtual interactions to meet with other core fans in person. These in-person meetings are usually tied to organizing fan events, such as planning promotional campaigns, or hosting fan gatherings. For example, core fans may come together offline to design and execute support projects for their idols, like billboard advertisements. However, when it comes to general fans, their engagement with core fans tends to remain online. Core fans often take deliberate steps to maintain their privacy, avoiding the disclosure of their real identities to general fans. This trend has also been noted in prior research on Chinese fan cultures (Xiao and Zheng, 2023), where it is highlighted that core fans carefully manage boundaries to preserve their influence and anonymity within the community. While some general fans might participate in offline events, they typically remain unaware of the true identities of the core fans who organize these activities.
Some core fans, according to our research, can even communicate directly with idols or their management teams. However, the precise nature of these relationships, such as the extent to which core fans influence idols’ decisions or public image, remains underexplored in current fandom studies. Core fans heavily depend on social media platforms to coordinate fan activities, disseminate information, and amplify their idols’ visibility. In some cases, core fans may gain access to privileged information through employees of these platforms, but this is neither a requirement for becoming a core fan nor a common trait among all core fans. Instead, the defining characteristics of core fans lie in their significant organizational skills and ability to rally large groups of fans within the community. These attributes grant core fans their elevated status and central role in fandom activities, regardless of whether they have insider access to information.
To clarify further, it is not the case that each core fan corresponds to a distinct and separate fan club. Some of the core fans we interviewed belong to the same fan club. Our research involved a total of 12 fan clubs. However, the core fans we interviewed in the same fan club all held different positions to ensure the diversity of the research. At the request of our participants, we have not disclosed with which fan clubs individual participants are affiliated, as doing so could enable some readers to potentially infer the specific fan clubs and potentially identify the real identities of the participants.
@@ 3.2. Participating Role
In fact, as researchers, we inevitably became part of the collective algorithmic action practices led by core fans, even shaping some aspects of these practices. For instance, during internal debates on how to optimize certain algorithmic practices, our feedback and observations, informed by academic perspectives, were occasionally incorporated into the core fans’ plans. This dual role of observer and inadvertent participant underscores the complex dynamic between researchers and the communities they study in ethnographic research, particularly in contexts where participation is integral to understanding collective actions.
@@
Furthermore, these two researchers also engaged in the algorithmic practices typical of general fans, dedicating 1 to 2 hours daily, from Monday to Friday, to perform routine activities such as repetitive liking, reposting, and commenting, and adjusting specific strategies across different platforms. The platforms where our researchers conducted these algorithmic practices included YouTube, Weibo, X (formerly Twitter), Douyin, TikTok, Xiaohongshu, and other popular platforms where fan activities are active. We did not participate in all 12 fan clubs’ algorithmic practices every day. Instead, based on the core fan clubs’ activity levels and relevance to our research questions, we selectively engaged with specific fan clubs on different days. Additionally, they actively joined and observed the online communities of general fans within the fan clubs corresponding to these 43 core fans, participating in their discussions.
We chose this participatory method as ethnographers because it provided a unique opportunity to gain in-depth insights into the nuanced, day-to-day dynamics of algorithmic collective actions that would be difficult to capture through interviews or passive observation alone. The participant-observer approach allowed us to understand the strategies, decision-making processes, and internal rhetoric in real time, which is crucial for examining the interplay between fan communities and algorithmic systems. This approach also fostered a level of trust and recognition from participants, enabling us to access sensitive and nuanced conversations that might otherwise remain hidden. This is particularly significant in the context of fan studies, where previous research has highlighted the vulnerability of fan communities and their guardedness towards outsiders (Brown, 2012). These dynamics make ethnography not only a valuable but often the primary method for studying such communities (Brown, 2012).
We acknowledge the need to reflect on how our involvement might have shaped the outcomes. By embedding ourselves within the community and actively participating in their practices, we were able to build rapport with participants, gaining their trust and, in turn, their willingness to share insights that are critical for understanding their algorithmic practices. This dual role of observer and participant aligns with the established practices in ethnographic fan research, where trust-building and immersive engagement are essential for generating authentic and comprehensive data.
3.3. Participants: Core Fans (N=43) and General Fans (N=15)
@@
The 43 Chinese core fans (CF1 to CF43) we interviewed all actively participated in organizing daily fan algorithmic actions. In addition, each organized over 30 specialized collective algorithmic actions targeting specific events, such as brand campaigns involving their idol, showcasing their extensive experience in coordinating such activities. The platforms impacted by the algorithmic actions discussed in these interviews included major fan activity platforms like YouTube, Weibo, X (formerly Twitter), Bilibili, Douyin, TikTok, Xiaohongshu, WeChat Channel and various music charts. Among the Chinese core fans we interviewed, 4 (CF1 to CF4) managed fan clubs with 100,000 to 200,000 followers on their social media accounts, 21 (CF5 to CF25) managed clubs with 200,000 to 500,000 followers, 11 (CF26 to CF36) managed clubs with 500,000 to 1,000,000 followers, and 7 (CF37 to CF43) managed clubs with more than 1,000,000 followers. In their view, following the idol might indicate an appreciation for the idol personally, but it does not necessarily imply active participation in algorithmic actions. In contrast, those who follow a fan club are much more likely to be highly motivated to engage in fan-driven algorithmic activities. Therefore, the data from these fan clubs can assist HCI researchers in understanding the scale of their organized algorithmic actions. These interviewed Chinese core fans could directly participate in and influence their fan clubs’ decisions and judgments regarding algorithmic folk understanding and collective algorithmic actions.
Based on the core fans we interviewed, they believe that about 10% of a fan club’s followers join in daily collective actions aimed at boosting their favorite celebrity’s visibility on social media. These daily actions are usually straightforward, such as liking, commenting, or sharing posts. However, when the tasks become more complicated, like organizing coordinated campaigns or using specific tactics to influence algorithms, only about 3% to 5% of followers participate. This number can vary depending on how loyal and united the general fans are. That said, during important events for their favorite celebrity, such as a new album release or a major award nomination, participation could go up to as much as 20%. It’s important to emphasize that these numbers are not precise but rather the subjective estimates shared by the core fans we spoke to.
Additionally, all the core fans interviewed have their own social media account besides their fan club, while each had over 5,000 followers on social media platforms, with a monthly average readership exceeding 2 million, and their social media content primarily focused on their idol and fan community, thus affirming their core fans status (sufficient influence driven by their own fan identity).
@@ 3.4. Data Collection
The timeline of the data collection process spanned three phases: online and on-site observations conducted from 2021 to 2023, semi-structured interviews conducted in 2022, and focus group sessions held in early 2023. This phased approach allowed the researchers to iteratively refine their understanding of fan-led algorithmic practices, ensuring a comprehensive exploration of the phenomenon.